OpenStudy (anonymous):

Would I use the equation y2-y1 m= ------ x2-x1 to find an equation of a line?

OpenStudy (anonymous):

yes. point slope formula

OpenStudy (anonymous):

to find the slope of the line.***

OpenStudy (anonymous):

so you would find "m" from y = mx+b

OpenStudy (anonymous):

Ohhhh so it would be y-y1 = m(x-x1) ?? Do you happen to know if there is a major difference between those two equations ( One in my question and teh one in this reply?)

OpenStudy (anonymous):

gotcha. Thank youuu :)

OpenStudy (anonymous):

they are the same equation

OpenStudy (anonymous):

divide both sides by (x-x1) and u get the other form

OpenStudy (anonymous):

Thats where I got confused :) Thank yous :)

OpenStudy (anonymous):

np